A train leaves the station traveling west at a constant rate of 45 mph. An express train leaves the same station 1 hour later heading west on the same route, traveling at a constant rate of 60 mph. How many hours will the first train have been traveling when the express train catches up to it?
Answer by TimothyLamb(4379)

---
head start trainA:
s = d / t
d = s * t
d = 45 * 1
d = 45 miles
---
relative speed:
60 - 45 = 15 mph
---
time to the meeting, after trainB departs:
t = d / s
t = 45/15
t = 3
---
answer:
trainA will have been traveling 4 hours when trainB catches up
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php